the diagonal of a square is 24 centimeter.how long is the side of the square/

Answer by nerdybill(7384)

the diagonal of a square is 24 centimeter.how long is the side of the square
.
apply Pythagorean theorem:
Let x = length of side of square
then
x^2 + x^2 = 24
2x^2 = 24
x^2 = 12
x = sqrt(12)
x = sqrt(2*2*3)
x = 2*sqrt(3) centimeters (exact answer)
.
Or, decimal approx:
x = 3.46 centimeters